### v1

@@ 6 Empirical Performance Results
Factory-provided system defaults were used in all cases. Default free-range unbound threads were used, with no pinning of threads to processors or NUMA nodes. We used the default version of the GNU g++ compiler as provided by the Ubuntu distributions.
We compared MCS, CLH, TWA, Tidex, Ticket Locks(TKT), HemLock, Reciprocating Locks(Recipro), Hapax, and Hapax augmented with the visible waiters optimization (HapaxVW).
@@
In Figure-1 we used a C++ benchmark that defines a simple structure type S that contains 5 32-bit int fields. The benchmark spawns concurrent threads, each of which loops, calling std::atomic<S>::exchange() to swap a local copy with a singleton global instance, followed by a non-critical section phase. We report the aggregate throughput rate, in completed exchange operations per second, at the end of a 10 second measurement interval, and we plot the median of 11 distinct runs in the figure. The number of threads is on the X-axis and aggregate throughput in the Y-axis. The C++ compiler and runtime implement std::atomic for such “large” objects by hashing the address of the instance into an array of pthread mutexes, and acquiring those as needed to implement the desired atomic action. Concurrent exchange operations on the same instance can thus result in contention on those underlying mutexes. We used the LD_PRELOAD facility to interpose on the pthread locking operations, replacing them with the algorithms of interest. Any required lock context is collocated adjacent to the fields in the lock body. In the non-critical phase, threads invoke a local xoroshiro128plus[Vigna-xoroshiro-PRNG] pseudo-random number generator (PRNG) to generate an integer uniformly distributed in [0,100) and then advance that same PRNG that many steps.
@@ 6.2 MutexBench benchmark
We repeated each data point 11 times, with independent runs, and reported the median throughput in Figure-2. For clarity and to convey the maximum amount of information to allow a comparison of the algorithms, the throughput on the -axis is offset from 0, and, as needed, the -axis is broken or log scale.
@@ 7 Conclusion
††margin: claim, novel, novelty, contribution ††margin: Lesson, Informs, Instructs, Guides ††margin: Exemplar; explainer ††margin: MRAT = most recently arrived thread ††margin: strikes compromise; trade-off; balance; tension ††margin: derangement, disorder ††margin: Pick random starting position and direction ††margin: incur; suffer; induce; generate; require ††margin: KPIUnlike CLH, and like TWA, no pointers shift thread ownership (using the sense of ownership employed by the Rust and C++ programming languages [ownership], where responsibility for disposal of the referent transfers with the pointer), which is appealing as no memory lifecycle issues can exist, making it easier to reason about memory-safety. In HemLock, Reciprocating Locks, MCS and MCSH, threads spin privately and locally on memory that remains owned by the waiting thread, but the address of that memory escapes, and is exposed to other threads, which can read and write that memory. Under Hapax Locks, no thread-local (thread owned) addresses leak, escape, and are exposed and shared for access by other threads.
@@ Acknowledgements.
We thank Peter Buhr at the University of Waterloo for access to ARMv8 and AMD systems. We also thank the anonymous PPoPP reviewers for their helpful comments.Appendix A Key Properties of Hapax Locks
@@ Appendix B Discussion
In managed runtime environments such as the HotSpot Java Virtual Machine, with balanced locking under the “synchronized” construct, such context information can be kept in the stack frame. In C++, context can be kept in std::lock_guard or std::scoped_lock instance, which is typically allocated on stack. Such constructions are an example of the the Resource Acquisition is Initialization(RAII)[RAII] idiom. Context can also be conveyed to the unlock operation via a variant API, which allows passing of information, by using a lambda-based interface, through extra fields in the lock body, or via data structures held in thread-local storage that track held locks.
@@
Hapax Locks are thread-oblivious[topc15-dice] in that one thread can acquire a lock, but a different thread, in possession of the associated context, communicated from the acquiring thread, can subsequently release the lock. As such, they can be used in NUMA-aware Cohort Locks, as bounded binary semaphores, or other use cases where thread-oblivious locks are necessary.
@@
Consider implementing try_lock in Tidex. We fetch and inspect the Arrive and Depart variables and find them equal, so the lock is apparently unlocked – apparently as we didn’t get an atomic snapshot of the two variables, as we used two loads. Ignoring the alternative ID issue, we then use CAS to try to swing Arrive to our thread’s identity. But the fields could have shifted underneath us, from other concurrent actions on the lock, yielding an ABA issue where our CAS was successful, but we then discover that Depart is changed, in which case we’d be forced to wait in try_lock. We might also try to revert and back-out our update of Arrive, with another CAS, but there is no guarantee of success as other arrivals may have updated Arrive. The same scenario can manifest in ticket locks, although it takes 4G roll-over of 32-bit Ticket and Grant fields to occur. If we are willing to pack and collocate Arrive and Depart fields (or Ticket and Grant) fields together, and then use a double-wide CAS, you can safely implement try_lock in that fashion. If we adapt ticket locks to use 64-bit fields, then, as roll-over and ABA are impossible, as a practical concern, we can safely implement try_lock. Hapax Locks also gives us a viable try-lock as we have 64-bits and non-recurring values. We further observe that try_lock is safe even with the Hapax Locks visible waiters and positive handover. During positive handover episodes (sustained contention), we skip updating the Depart field (we resume updating Depart when contention subsequently abates) but we also know that Arrive and Depart must be different during such an episode, so the lock appears in locked state, and any try_lock attempt will fail.
@@
Hapax Locks are trivially amenable to using more sophisticated waiting techniques on the slots, such as WFE on ARM, MONITOR-MWAIT on Intel, or futex-based waiting.
@@ Appendix F Additional discussion regarding the Visible Waiters variant
We observe, however, that Hapax Locks variant in Listing-LABEL:Listing:Nonce-Visible-Waiters could be made threading-model aware, if, when waiting, immediately before they dismount from and abdicate a kernel thread (switching that kernel thread to carry another runnable 777A runnable thread is ready and eligible for dispatch, but is not currently running on kernel thread user-mode virtual) they also surrrender their slot in the waiting array in order to be altruistic and “polite” and decrease the odds of collisions in the array for other running threads. When the virtual thread is eventually dispatched and re-mounted on a kernel thread, it would then recheck Depart, and, if necessary, try to reclaim the slot and reinstall itself in the array.

### v2

@@ 6 Empirical Performance Results
Factory-provided system defaults were used in all cases. Unless otherwise noted, default free-range unbound threads were used, with no pinning of threads to processors or NUMA nodes. We used the default version of the GNU g++ compiler as provided by the Ubuntu distributions.
We compared MCS, CLH, TWA, Tidex, Ticket Locks, HemLock, Reciprocating Locks(Recipro), Hapax (HapaxIW), and Hapax augmented with the visible waiters optimization (HapaxVW).
@@
In Figure-1 we used a C++ benchmark, running on the ARMv8 system, that defines a simple structure type S that contains 5 32-bit int fields. The benchmark spawns concurrent threads, each of which loops, calling std::atomic<S>::exchange() to swap a local copy with a singleton global instance, followed by a non-critical section phase. We report the aggregate throughput rate, in completed exchange operations per second, at the end of a 10 second measurement interval, and we plot the median of 11 distinct runs in the figure. The number of threads is on the X-axis and aggregate throughput in the Y-axis. The C++ compiler and runtime implement std::atomic for such “large” objects by hashing the address of the instance into an array of pthread mutexes, and acquiring those as needed to implement the desired atomic action. Concurrent exchange operations on the same instance can thus result in contention on those underlying mutexes. We used the LD_PRELOAD facility to interpose on the pthread locking operations, replacing them with the algorithms of interest. Interposition allows us to change the underlying mutex without changing the application binary, by simply changing the LD_PRELOAD environment variable to reference various lock implementation libraries, which are loaded at process startup time. This allows for a fair comparision, but, like all pthreads implementations, precludes any compile-time optimizations that might otherwise be enabled at locking sites, between the application code and the locking algorithm. Any required lock context is collocated adjacent to the fields in the lock body.
In the non-critical phase, threads invoke a local xoroshiro128plus[Vigna-xoroshiro-PRNG] pseudo-random number generator (PRNG) to generate an integer uniformly distributed in [0,100) and then advance that same PRNG that many steps.
6.2 LevelDB
In Figure-2 we used the “readrandom” benchmark in LevelDB version 1.23
database888leveldb.org, again on the ARMv8 system, varying the number of
threads and reporting throughput
from the median of 5 runs of 50 seconds each.
Each thread loops, generating random keys and then tries to read the associated value from
the database.
We first populated a database999db_bench —-threads=1
––benchmarks=fillseq —-db=/tmp/db/
and then collected data101010db_bench —-threads=threads
—-benchmarks=readrandom
—-use_existing_db=1
—-db=/tmp/db/ —-duration=50.
We made a slight modification to the db_bench benchmarking
harness to allow runs with a fixed duration that reported aggregate throughput.
LevelDB uses coarse-grained locking, protecting the database with a single central mutex:
DBImpl::Mutex. Profiling indicates contention on that lock via leveldb::DBImpl::Get().
The results in Figure-2 largely echo the moderate contention results in Figure-3.
@@ 6.3 MutexBench benchmark
We repeated each data point 11 times, with independent runs, and reported the median throughput in Figure-3. For clarity and to convey the maximum amount of information to allow a comparison of the algorithms, the throughput on the -axis is offset from 0, and, as needed, the -axis is broken or log scale.
@@
AMD-N1 reflects a configuration of the AMD system where constrained execution to just one NUMA node, with 256 processors (instead of the full complement of 512), in order to better examine NUMA performance effects111111numactl –cpunodebind 1 –membind 1.
@@ 7 Conclusion
††margin: claim, novel, novelty, contribution ††margin: Lesson, Informs, Instructs, Guides ††margin: Exemplar; explainer ††margin: MRAT = most recently arrived thread ††margin: strikes compromise; trade-off; balance; tension ††margin: derangement, disorder ††margin: Pick random starting position and direction ††margin: incur; suffer; induce; generate; require ††margin: KPIUnlike CLH, and like TWA, no pointers shift thread ownership in Hapax Locks (using the sense of ownership employed by the Rust and C++ programming languages [ownership], where responsibility for disposal of the referent transfers with the pointer), which is appealing as no memory lifecycle issues can exist, making it easier to reason about memory-safety. In HemLock, Reciprocating Locks, MCS and MCSH, threads spin privately and locally on memory that remains owned by the waiting thread, but the address of that memory escapes, and is exposed to other threads, which can read and write that memory. Under Hapax Locks, no thread-local (thread owned) addresses leak, escape, and are exposed and shared for access by other threads.
@@ Acknowledgements.
We thank Peter Buhr at the University of Waterloo for access to ARMv8 and AMD systems. We also thank the anonymous PPoPP reviewers for their helpful comments.References
Appendix A Key Properties of Hapax Locks
@@ Appendix B Discussion
We observe collisions in the wait array can be either intra-lock or inter-thread.
Hapax Locks with invisible waiters can allow threads in different processes to safely synchronize on shared inter-process locks, as long as a shared singleton nonce allocator is used, and the waiting array is mapped into each address space. This is enabled by virtue of the fact that no addresses are shared between participating threads.
@@
In managed runtime environments such as the HotSpot Java Virtual Machine, with balanced locking under the “synchronized” construct, such context information can be kept in the stack frame. In C++, context can be kept in std::lock_guard or std::scoped_lock instance, which is typically allocated on stack. Rust’s Mutex<T> is similar. Such constructions are an example of the the Resource Acquisition is Initialization(RAII)[RAII] idiom. Context can also be conveyed to the unlock operation via a variant API, which allows passing of information, by using a lambda-based interface, through extra fields in the lock body, or via data structures held in thread-local storage that track held locks.
@@
Hapax Locks are thread-oblivious[topc15-dice] in that one thread can acquire a lock, but a different thread, in possession of the associated context, communicated from the acquiring thread, can subsequently release the lock. As such, they can be used in NUMA-aware Cohort Locks, as bounded binary semaphores, or other use cases where thread-oblivious locks are necessary. In addition, with the uptake of asynchronous futures, it is more common to find situations where one thread accquires a lock, and some other thread may subsequently release that lock 121212forexample:https://github.com/serenedb/serenedb/blob/main/libs/basics/unshackled_mutex.h.
@@
Consider implementing trylock in Tidex. We fetch and inspect the Arrive and Depart variables and find them equal, so the lock is apparently unlocked – apparently as we didn’t get an atomic snapshot of the two variables, as we used two loads. Ignoring the alternative ID issue, we then use CAS to try to swing Arrive to our thread’s identity. But the fields could have shifted underneath us, from other concurrent actions on the lock, yielding an ABA issue where our CAS was successful, but we then discover that Depart is changed, in which case we’d be forced to wait in trylock. We might also try to revert and back-out our update of Arrive, with another CAS, but there is no guarantee of success as other arrivals may have updated Arrive. The same scenario can manifest in ticket locks, although it takes 4G roll-over of 32-bit Ticket and Grant fields to occur. If we are willing to pack and collocate Arrive and Depart fields (or Ticket and Grant) fields together, and then use a double-wide CAS, you can safely implement trylock in that fashion. If we adapt ticket locks to use 64-bit fields, then, as roll-over and ABA are impossible, as a practical concern, we can safely implement trylock, where the operator first fetches the Grant field, observing value , and then tries to optimistically apply atomic compare-and-swap to swing the Ticket field from to .
Hapax Locks also gives us a viable trylock as we have 64-bits and non-recurring values. We further observe that trylock is safe even with the Hapax Locks visible waiters and positive handover. During positive handover episodes (sustained contention), we skip updating the Depart field (we resume updating Depart when contention subsequently abates) but we also know that Arrive and Depart must be different during such an episode, so the lock appears in locked state, and any trylock attempt will fail.
@@
Hapax Locks are trivially amenable to using more sophisticated waiting techniques on the slots, such as WFE on ARMv8, MONITOR-MWAIT on Intel, or futex-based waiting. On ARMv8 systems we have observed that waiting via WFE makes a signficant improvement in contented throughput, compared to YIELD, for locks that uses private spinning, but that UMONITOR and MONITORX made no appreciable improvement for throughput on AMD and Intel processors.
@@
Hapax values can also be used in fashion similar to seqlocks[seqlock1, seqlock2]. Each record being protected by a hapax-seqlock has a payload area, a WriteFirst and WriteLast field, containing hapax values. To write the record, a thread creates a hapax value, stores that value into WriteFirst, then updates the payload, and finally stores the same hapax value into WriteLast. Note that writers, even when there are concurrent writers, never block or wait. Concurrent readers will first fetch WriteLast, fetch a tentative copy of the payload, and then fetch WriteFirst. Critically, readers fetch the hapax fields in reverse order relative to the write order. If the two fetched values match, then the reader has confirmed that the payload fetched is a consistent atomic snapshot, and is safe to use. Otherwise the snapshot is inconsisent, and the reader takes action to respond accordingly. The snapshot can be inconsistent because at least one concurrent write is in progress, or when previous multiple concurrent writers overlapped and interleaved accesses, leaving the record in an inconsistent state. This technique is useful for concurrent caches, for instance, where inconsisent values can be reconstituted and re-installed, recovering from the inconsistent state. We expect the payload to be trivially copyable, overwritable without risk of leaks, and to not require destructors, as the value can be “stomped” by concurrent writers. We observe that using this technique in modern C++ requires particular care[SeqLockMemcpy].
@@
We call out a confounding benchmark phenomenon we call the re-arrival effect. We say a lock algorithm is unlock-terse if, under sustained contention, it avoids updating lock body fields in unlock, and unlock-prolix if it usually updates the lock body in unlock. Unlock-prolix lock algorithms that always write into the shared lock body, in unlock, such was Ticket Locks, will usually suffer more coherence traffic than do unlock-terse algorithms – such as MCS and CLH – that avoid, under sustained contention, updating the global lock body state in unlock.
Say we have a maximum contention benchmark with an empty non-critical sections, and sustained contention. Threads release the lock in unlock and then immediately re-arrive in lock to re-acquire the same lock. But, with an empty critical section, and assuming all the lock body fields are collocated on the same cache line, the updates by an unlock-prolix algorithm in unlock will drive the line(s) underlying the lock body into M-state, and the subsequent lock operation will re-access those same lines while they still reside in the local cache in M-state, as no other lock operations have had an opportunity to arrive or depart and invalidate the line. This effect somewhat distorts performance and confers a relative performance benefit to prolix lock algorithms in the case of maxmimum sustained contention with no intervening non-critical section. The degree of impact is specific to the architecture and the lock algoirthm – whether prolix or not.
The lock operator, for an unlock-prolix algorithm, can avoid coherence misses that it might otherwise incur if threads arrived between the lock and unlock.
We note that the existence of lock context passed through the lock body can trigger the re-arrival effect, even if the algorithm is normally unlock-terse.
Using randomized non-critical section durations acts to mitigate this effect, and is also arguably more faithful to the real-world application behavior we are trying to model in benchmarks.
Similar to CLH and Reciprocating locks, no explicit queue of waiting threads is made manifest. Waiting threads know only the identity of their immediate successors. As such, these locks can not trivialy be adapted, when implemented within a pthread mutex-condvar environment, to enable the classic wait morphing optimization, where signaling a condition variable within a critical section simply moves threads from the condition variable’s wait set to the mutex’s queue of waiting threads.
Another benchmarking pitfall is as follows. Say we have a FIFO lock algorithm, and, for the purposes of this explanation, an empty critical section. It it then common to find a long-standing repeating cyclic admission schedule. Often, this schedule is set via happenstance at startup, but then persists for long periods. A given schedule can be either favorable or unfavorable, for throughput, in terms of the system topology. Using NUMA as an example, a given schedule might reflect either many or few lock migrations, which in turn will reflect in the throughput achieved by that schedule. But other topology structures, such as AMD’s CCX clusters, can yield similar favoratism. This, in turn, can result in higher run-to-run variance, if we assume the initial schedule arrises in a mostly random fashion. This effect can be mitigated in part by injecting randomized non-critical section delays, in order to shuffle or shift the schedule and avoid long-standing persistent schedules.
@@ Appendix F Additional discussion regarding the Visible Waiters variant
We observe, however, that Hapax Locks variant in Listing-LABEL:Listing:Nonce-Visible-Waiters could be made threading-model aware, if, when waiting, immediately before they dismount from and abdicate a kernel thread (switching that kernel thread to carry another runnable 131313A runnable thread is ready and eligible for dispatch, but is not currently running on kernel thread user-mode virtual) they also surrrender their slot in the waiting array in order to be altruistic and “polite” and decrease the odds of collisions in the array for other running threads. When the virtual thread is eventually dispatched and re-mounted on a kernel thread, it would then recheck Depart, and, if necessary, try to reclaim the slot and reinstall itself in the array.